OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

jimthompson5910 (jim_thompson5910):

first distribute to get 3(x + 1) = −2(x − 1) + 6 3x + 3*1 = −2x - 2(-1) + 6 3x + 3 = −2x + 2 + 6

jimthompson5910 (jim_thompson5910):

then combine like terms 2 and 6 on the right side to get 3x + 3 = -2x + 8 what's your next step?

OpenStudy (anonymous):

I don't know, lol. I really need help I am bad at algebra, and need the answer. 1, 4, 5, 25 are my choices..

jimthompson5910 (jim_thompson5910):

you want to isolate x

jimthompson5910 (jim_thompson5910):

so you need to get all the x terms to one side and everything else to the other

jimthompson5910 (jim_thompson5910):

so you need to move that -2x over you do this by adding 2x to both sides to get 3x + 3 + 2x = 8 then you subtract 3 from both sides to get 3x + 2x = 8 - 3 now combine like terms 5x = 5 finally divide both sides by 5 to isolate x fully x = 5/5 x = 1

OpenStudy (anonymous):

Lmao thanksss(:

OpenStudy (anonymous):

Oops, disregard my answer

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

can you answer all the questions im about to put here?

jimthompson5910 (jim_thompson5910):

I'll do a few, but not all

OpenStudy (anonymous):

Ighttt.

OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? 576 240 144 960

jimthompson5910 (jim_thompson5910):

first subtract 0.6x from both sides (you want to move it over to the other side) x = 0.6x + 384 x - 0.6x = 384 0.4x = 384 what's your next step?

OpenStudy (anonymous):

Divide .4x on both sides...i think

jimthompson5910 (jim_thompson5910):

close, divide both sides by 0.4 to isolate x

jimthompson5910 (jim_thompson5910):

0.4x = 384 x = 384/0.4 x = ???

OpenStudy (anonymous):

960?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

hell ya. lol. uhmm now this one.

OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

jimthompson5910 (jim_thompson5910):

−2(x + 3) = −2(x + 1) − 4 −2x - 6 = −2x - 2 − 4... distribute -2x - 6 = -2x - 6 ... combine like terms. see how to finish up?

OpenStudy (anonymous):

uhh...cancel out 6..?

jimthompson5910 (jim_thompson5910):

good, add 6 to both sides -2x - 6 = -2x - 6 -2x - 6+6 = -2x - 6+6 -2x + 0 = -2x + 0 -2x = -2x what's next?

OpenStudy (anonymous):

2 3 All real numbers No solution these are my options

OpenStudy (anonymous):

divide two.?

jimthompson5910 (jim_thompson5910):

yep divide both sides by -2 -2x = -2x x = x this is ALWAYS true for ANY real number

jimthompson5910 (jim_thompson5910):

so that's why the solution set is the set of all real numbers

OpenStudy (anonymous):

the answer is all real numbers?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

pick any number you want, it will work for x when you plug it in

OpenStudy (anonymous):

thankssss dawwgg lol

OpenStudy (anonymous):

Solve for x: 3x − 5 = 2x + 6

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

last one

jimthompson5910 (jim_thompson5910):

ok try to work it out and show me how far you get

OpenStudy (anonymous):

3x − 5 = 2x + 6 im confused on the first step.

OpenStudy (anonymous):

How am i supposed to cancel out 5 and 6 if they arent the same numbers?

jimthompson5910 (jim_thompson5910):

ok you want to move all the x terms to one side, everything else to the other

jimthompson5910 (jim_thompson5910):

how do you move the 2x over?

OpenStudy (anonymous):

idk what to doo. god i hate algebra.

jimthompson5910 (jim_thompson5910):

on the right side, you're adding 2x undo this by subtracting 2x

jimthompson5910 (jim_thompson5910):

so that's why you subtract 2x from both sides

jimthompson5910 (jim_thompson5910):

3x - 5 = 2x + 6 3x - 5 - 2x = 6 how do you move that -5 over?

OpenStudy (anonymous):

I dont know what im talking aboutt lol. can you please tell me the answer? i suck at this pellet.

jimthompson5910 (jim_thompson5910):

its ok not to know, but don't give up

jimthompson5910 (jim_thompson5910):

you don't suck, you just need practice, that's all

OpenStudy (anonymous):

Why does it say pellet? lmao i typed s-h-i-t

OpenStudy (anonymous):

Can you please tell me?

jimthompson5910 (jim_thompson5910):

it's the censor/filter lol

jimthompson5910 (jim_thompson5910):

I can tell you, but you have to try a bit first

OpenStudy (anonymous):

Goddddddd. this bullpellet is so fluttering annoyinggg.

jimthompson5910 (jim_thompson5910):

this will be on a test someday and you'll have to know how to solve equations

OpenStudy (anonymous):

I am taking a test right now lol xD

jimthompson5910 (jim_thompson5910):

oh...

OpenStudy (anonymous):

Ya...So...uh..

jimthompson5910 (jim_thompson5910):

sry can't help, but if it weren't a test and you had more time, then I'd love to I can see why you're so stressed now. I hate tests too

OpenStudy (anonymous):

I figured it outt! its 11, lol. I used my calculator and put every choice in for x and got the same answer. yaa.

jimthompson5910 (jim_thompson5910):

nice work, see I knew you could do it

OpenStudy (anonymous):

Jim your a champion, very patient, understanding and helpful, keep up the good work